Ask
your own question, for FREE!
Mathematics
11 Online
Diana has a box containing 6 blue, 8 red, 5 purple, 9 green, and 2 clear marbles that are all the same size and shape. What is the probability of randomly choosing a clear marble on the first pick, replacing it, and then randomly choosing a purple marble on the second pick? A- 1/6 B- 1/15 C- 1/87 D- 1/90
Still Need Help?
Join the QuestionCove community and study together with friends!
since these are independent events (and you place the first marble back), the answer is : 1. P(drawing a clear) * P(drawing a purple) 2. there are 30 marbles, and 2 clear marbles, the probability of drawing a clear is 2/30 3. once you put one back, there are still 30 marbles, and 5 purple marbles, and the probability of drawing a purple is 5/30 4. so the answer is 2/30 * 5/30 = 10/900 = 1/90
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
11 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal